OpenStudy (osanseviero):

I have a circle. I know 2 points that are in the circunference. I also know one line that touches the center. How can I calculate the center?

OpenStudy (osanseviero):

What I thought would be: Getting the equation of the line of the two points, then the perpendicular, then saying it is equal to the other line.

OpenStudy (anonymous):

 find slope of line AB,m=(y2-y1)/(x2-x1) slope of line PD=-1/m=M find mid-point L and eq. of line through L by y-y3=M(x-x3) find the point of intersection of this line PD with the given line say MN. it is centre.

OpenStudy (osanseviero):

I forgot about midpoing D: D: D: